# Magnetic field generation in mergers of massive main sequence stars

### v1

@@
Magnetic fields are found in many astrophysical objects, ranging from galaxy clusters to the interstellar medium of galaxies and magnetars. Strong magnetic fields are also observed in massive stars, but it is still unclear how they are generated. There are different theories for their origin: the magnetic fields could be fossil fields from star formation, they could be generated in dynamo processes during stellar evolution, or they could be generated during a merger of two stars. Here, we show how magnetic fields are generated in simulations of mergers of massive main-sequence stars, conducted with the 3D magnetohydrodynamics code Arepo. In these simulations, a and a main sequence star merge, whereby the more massive star transfers mass to the less massive one until it is finally disrupted, forming a core–disk structure. During the initial mass transfer, magnetic fields start to be amplified in the accretion stream, saturating after the disruption at an equipartition level. Correspondingly, the magnetic energy starts being present at small scales, growing to larger scales during the merger. During the merger, the magnetic pressure is much smaller than the gas pressure, thus the dynamical impact of the magnetic fields on the merger is small. After the merger, both are similar in some regions and the magnetic field shows a large-scale structure that is mostly toroidal and that is expected to be stable. Thus, our simulations show that magnetic fields in massive stars can originate from a merger of two main sequence stars.
@@
hydrodynamics – methods: numerical – stars: general1 Introduction
@@
The merger of two stars is a common phenomenon in stellar astrophysics and may invoke compact objects such as neutron stars and black holes, but also stars during their main nuclear-burning cycles. While the former are observed, e.g., in gravitational-wave events (Abbott et al. 2016, 2017), the latter lead to outbursts that are called luminous red novae with V1309 Sco and V838 Mon being prototypical events (Munari et al. 2002; Tylenda et al. 2011). Merged stars may be rejuvenated (e.g. Hellings 1983; Braun & Langer 1995; Dray & Tout 2007; Schneider et al. 2016) and can appear as blue stragglers in star clusters and stellar associations (e.g. Hills & Day 1976). Some of the most massive stars known might in fact be merger remnants (Schneider et al. 2014) and also the Great Eruption of Car in the 1840s that created the Homunculus Nebula has been suggested to be the result of a merger event (e.g. Smith et al. 2018; Hirai et al. 2021). Some merged stars may also give rise to unusual supernova progenitors such as the surprising blue supergiant that lead to SN 1987A (Podsiadlowski & Joss 1989; Podsiadlowski et al. 1990, 1992; Morris & Podsiadlowski 2007; Menon & Heger 2017) while others are thought to be the progenitors of interacting and even superluminous supernovae (Justham et al. 2014). Still, the merging phase itself and also the evolution and final fates of merged stars leave many open questions. Schneider et al. (2025) present a general introduction to stellar mergers and the related common-envelope phase.
@@
Mergers can be induced in various ways: by dynamical encounters in dense stellar systems such as star cluster (cf. “head-on collisions”), by orbital decay of binaries in dense interstellar/circumbinary media such as in AGN disks or during the formation of stars, by orbital dynamics in triple-star systems and other multiples (e.g. von Zeipel–Lidov–Kozai mechanism), and by mass transfer in isolated binary stars. The latter channel is supposed to lead to mergers for about 25% of all massive stars (; Sana et al. 2012) such that – when integrating over a constant star-formation rate in a stellar population like in the Milky Way – about 10% of all massive stars are expected to be remnants of mergers from isolated binary evolution (de Mink et al. 2014). Indeed, the binary evolution computations of Henneco et al. (2024) show that at least 12% of massive, mass-transferring binaries will merge, and this fraction goes up to if fully conservative mass transfer is assumed (Henneco 2025).
@@
Ever since the discovery of surface magnetic fields in the A-type star 78 Vir by Babcock (1947), the origin of strong surface magnetism in massive OBA stars is debated. It has long been speculated that strong and large-scale magnetic fields are produced in stellar mergers and that they can persist over the long evolutionary timescales of stars (see e.g. Ferrario et al. 2009; Wickramasinghe et al. 2014). Today, direct simulations of essentialy every flavor of stellar mergers and other dynamic stellar phases show the efficient amplification of magnetic fields. This is the case in mergers of binary white dwarfs (e.g. Zhu et al. 2015; Pakmor et al. 2024), white dwarfs and neutron stars (e.g. Morán-Fraile et al. 2024), binary neutron stars (e.g. Kiuchi et al. 2024), low- and high-mass main-sequence (MS) stars (e.g. Schneider et al. 2019, hereafter Paper I; Ryu et al. 2024; Vynatheya et al. 2025) and also common-envelope events (e.g. Ohlmann et al. 2016b; Ondratschek et al. 2022; Vetter et al. 2024; Gagnier & Pejcha 2024).
@@ 2.2 Hydrodynamics simulations
For the merger simulations, we need to create initial conditions that consist of two stars that are about to merge. For these single stars, we generate stable models by following the method outlined in Ohlmann et al. (2017). We compute the stellar structure with Mesa and map it to a grid with cell centers on spherical shells distributed according to a HEALPix tesselation. Then, we run a relaxation because slight mismatches between pressure gradients and gravitional forces can induce spurious velocities that need to be damped away. We carry out the relaxation over ten dynamical timescales as given by the sound crossing time to ensure that the models remain stable. We give some parameters of the two initial stellar models in Table 1.
@@
where is a timescale on which the components merge and is the index of the cell. For this step, we take s and run the simulation for about 1.5 orbits with the loss term and continue without it from there on. For comparison, we start one model without the loss term after only half this time and it shows a very similar evolution. The simulations use 3.4 to 4 million cells in the high resolution runs and about 0.43 to 0.45 million cells in the low resolution runs. We set up the initial magnetic field as a dipole field, similar as in Ohlmann et al. (2016b), with a surface field of 1 G. We run one additional simulation without magnetic fields for comparison. We provide an overview of the models in Table 2.
@@ 3.2 Evolution of the merger
The morphology of the flow during the merger is depicted in Fig. 3, showing slices of the density through the orbital plane and perpendicular to it, and in Fig. 4, showing slices of the passive scalar tracing material of the primary star through the orbital plane. Initially, mass transfer from the primary star to the secondary star sets in via the inner Lagrange point L1. This stream hits the secondary and leads to shear flows near the surface of the secondary. In addition, mass and also angular momentum are lost through the outer Lagrange points L2 and L3 (see, e.g., Fig. 3 at ), thereby shrinking the orbit. The orbital shrinkage accelerates until the stars merge: the initially more massive primary star is disrupted and distributed around the secondary star (see Figs. 3 and 4 at ). Directly after the merger, the cores of the stars are still separate (see central part of Fig. 4 at ). Shear in this central part causes hydrodynamical instabilities, mixing the cores and resulting in a symmetric merger product (Figs. 3 and 4 at ). Due to the large angular momentum of the binary, a disk-like structure forms around the rotating core with a mass of about (see Fig. 3 at in the – plane). The central core mainly originates from the less massive secondary, while the fraction of material from the primary increases outwards. The post-merger structure is described in more detail in Sect. 3.4.
The simulations of the merger start after the orbit of the binary system has shrunk in an additional step to (Model 1-3) or (Model 4), corresponding to orbital periods of about d and d, respectively. Strong mass transfer (about ) develops from the more massive primary star onto the secondary. The orbital separation decreases until the stars merge after roughly 23 (Model 1) to 35 (Model 4) orbits111For larger initial separations, the system becomes even more stable and takes more orbits to merge, rendering the simulations computationally unfeasible. (Fig. 2). As expected, Model 4 with a larger initial separation takes longer to merge than Models 1–3 with a smaller initial separation. Here, we define as the point in time where the orbital shrinkage becomes large, i.e., where the corresponding time scale becomes small and fulfills (where is the semi-major axis). Fig. 2 also shows that the distance varies with the orbital motions, i.e., the orbit is slightly eccentric. This is probably caused by the system not being perfectly Keplerian after the first step of artificially shrinking the system. The effect is smaller for the high-resolution Model 1.
@@ 3.3 Magnetic field amplification
The amplification continues in the mass stream from the primary to the secondary as it creates shear flows outside the Roche lobes of both stars (Fig. 6 at d). At this stage, we find that the flow is unstable to the magneto-rotational instability (MRI, Balbus & Hawley 1998). We estimate the length and time scale of the fastest growing MRI channel in our simulations (cf. Rembiasz et al. 2016): this yields a timescale of about 0.5 d and a wavelength of about which is similar to the scales in our simulation (time scale: 0.2–1 d, length scale:). Moreover, the strong amplification of the magnetic field starts as soon as the wavelength of the fastest growing MRI channel is resolved by a few hundred cells.
@@
The amplification is slowing down and reaches saturation at later times, when the ratio of magnetic to turbulent kinetic energy approaches the percent level and comes closer to equipartition (Fig. 5, middle panel). As a proxy for the turbulent kinetic energy, we use the kinetic energy in the radial and direction, , which is not affected by the large-scale angular motion (see the comparison in Fig. 5, bottom panel).
The length scales can also be identified in power spectra of the kinetic and magnetic energy (Fig. 7). Most of the kinetic energy resides at large scales and low wave numbers for all times, as expected for the rotation of the stellar cores and of the post-merger product (Fig. 7, upper panel). The power spectrum follows a power law to smaller scales (exponent between and ), probably due to turbulence seeded by flow instabilities. The exponents are smaller in magnitude than the Kolmogorov exponent of that would be expected for a homogenous, isotropic turbulent cascade. Plots of power spectra of the magnetic energy (Fig. 7, lower panel), however, show that magnetic energy is injected at small scales (about –) and propagates to larger scales during the amplification (see d to d). After the cores have merged, redistribution of the magnetic energy leads to a similar spectrum at 5 d as for the kinetic energy, with most of the energy at large scales (about –). The exponents of the power law have been obtained by a fit to for between 0.1 and and are 1.38, , and for d, d, and d, respectively. This behavior is similar to the MRI simulations by Bhat et al. (2016), in which large-scale modes only grow late in the evolution when the amplification nears saturation. As mentioned there, large-scale magnetic fields seem to appear when the amplification process saturates (see also Ebrahimi et al. 2009).
@@ 3.4 Structure of the merger product
The large angular momentum in the binary system leads to the formation of a disk around a rotating core in the merger product (compare Fig. 3 at d). The density profile shows strong asymmetries between polar and equatorial regions indicating that the boundary between core and disk is located roughly at (Fig. 8). Moreover, the mass profile increases only in the equatorial region for radii above (Fig. 8, lower panel). Thus, almost all mass for radii above is contained in the equatorial region, i.e., in a disk-like structure. At d, the core contains a mass of about , and the disk a mass of about . Moreover, the disk comprises 58% of the total angular momentum.
@@
The velocity at the core–disk boundary is about 700 km s-1 at d, i.e. 76% of the critical Keplerian rotational velocity. As can be seen in Fig. 9, second panel, the rotation curve in the disk always stays sub-Keplerian. After six days, this disk has completed about 170 rotations at the base and about 0.3 rotations at the edge. The ratio of kinetic energy to gravitational energy decreases from 0.07 at d to 0.05 at d and stays well below the limit of 0.14 for non-axisymmetric instabilities in rotating fluids (Ostriker & Tassoul 1969; Chandrasekhar 1970; Friedman & Schutz 1975; Lai & Shapiro 1995). Thus, no secular or dynamical, non-axisymmetric instabilities of the configuration are expected.
@@
As already seen during the evolution of the merger (cf. Fig. 4), the very central part of the merger product originates from the secondary star, whereas the primary star is disrupted around the core. Only the inner of the product are dominated by material from the secondary, the rest of the core stems to similar parts from both stars. The disk, however, is predominantly formed from material of the primary star. This distribution is expected from the entropy and thus buoyancy profiles of the progenitor stars: due to the similar mass and hence also similar evolutionary stage, the entropy is similar and mixing is expected over large parts of the profile. Because the secondary has a lower mass, its entropy is slightly smaller and thus its material dominates the central part. This behaviour is similar to the case of direct collisions as explored by Glebbeek & Pols (2008) and Glebbeek et al. (2013), especially their case S.
@@
The magnetic field structure of the final merger product is visualized in Fig. 10. The field structure is mostly toroidal; the energy of the toroidal magnetic field component is about 80-85% of the total magnetic energy. In this range, magnetic field configurations are expected to be stable in stars (Braithwaite & Spruit 2004; Braithwaite & Nordlund 2006a). Moreover, the configuration remains stable after the final merger and this time already corresponds to a few Alfvén timescales. However, in the regime of fast rotation (), the evolutionary timescale of the magnetic field is , where is the Alfvén timescale (Braithwaite & Cantiello 2013). Hence, great changes in the magnetic field are not expected to be seen over the timescales simulated here after the field has reached an equilibrium configuration. We expect Ohmic decay to operate on timescales that are comparable to the lifetime of the star (for more details, see Paper I). Hence, the magnetic field of the star should remain present at least until it reaches the main sequence. Modeling the subsequent stellar evolution in more detail shows that the merger product could evolve into a MS star with strong magnetic fields, similar to Sco (Paper I; Schneider et al. 2020).

### v2

@@
Magnetic fields are found in many astrophysical objects, ranging from galaxy clusters to the interstellar medium of galaxies and neutron stars. Strong magnetic fields are also observed in massive stars, but it is still unclear how they are generated. There are different theories for their origin: the magnetic fields could be fossil fields from star formation, they could be generated in dynamo processes during stellar evolution, or they could be generated during a merger of two stars. Here, we show how magnetic fields are generated in simulations of mergers of massive main-sequence stars, conducted with the 3D magnetohydrodynamics code Arepo. In these simulations, a and a main sequence star merge, whereby the more massive star transfers mass to the less massive one until it is finally disrupted, forming a core–disk structure. During the initial mass transfer, magnetic fields start to be amplified in the accretion stream, saturating after the disruption at an equipartition level. Correspondingly, the magnetic energy starts being present at small scales, growing to larger scales during the merger. During the merger, the magnetic pressure is much smaller than the gas pressure, thus the dynamical impact of the magnetic fields on the merger is small. After the merger, both are similar in some regions, and the magnetic field shows a large-scale structure that is mostly toroidal and that is expected to be stable. Thus, our simulations show that magnetic fields in massive stars can originate from a merger of two main sequence stars.
@@
Magnetohydrodynamics – Methods: numerical – binaries: general – Stars: magnetic field – Stars: massive1 Introduction
@@
The merger of two stars is a common phenomenon in stellar astrophysics and may invoke compact objects such as neutron stars and black holes, but also stars during their main nuclear-burning cycles. While the former are observed, e.g., in gravitational-wave events (Abbott et al. 2016, 2017), the latter lead to outbursts that are called luminous red novae with V1309 Sco and V838 Mon being prototypical events (Munari et al. 2002; Tylenda et al. 2011). Merged stars may be rejuvenated (e.g. Hellings 1983; Braun & Langer 1995; Dray & Tout 2007; Schneider et al. 2016) and can appear as blue stragglers in star clusters and stellar associations (e.g. Hills & Day 1976). Some of the most massive stars known might in fact be merger remnants (Schneider et al. 2014) and also the Great Eruption of Car in the 1840s that created the Homunculus Nebula has been suggested to be the result of a merger event (e.g. Smith et al. 2018; Hirai et al. 2021). Some merged stars may also give rise to unusual supernova progenitors such as the surprising blue supergiant that lead to SN 1987A (Podsiadlowski & Joss 1989; Podsiadlowski et al. 1990, 1992; Morris & Podsiadlowski 2007; Menon & Heger 2017) while others are thought to be the progenitors of interacting and even superluminous supernovae (Justham et al. 2014). Still, the merging phase itself and also the evolution and final fates of merged stars leave many open questions. Schneider et al. (2025) present a general introduction to stellar mergers and the related common-envelope phase, while Schneider (2025) reviews stellar mergers in depth.
@@
Mergers can be induced in various ways (cf. Schneider 2025): by dynamical encounters in dense stellar systems such as star cluster (cf. “head-on collisions”), by orbital decay of binaries in dense interstellar/circumbinary media such as in AGN disks or during the formation of stars, by orbital dynamics in triple-star systems and other multiples (e.g. von Zeipel–Lidov–Kozai mechanism), and by mass transfer in isolated binary stars. The latter channel is supposed to lead to mergers for about 25% of all massive stars (; Sana et al. 2012) such that – when integrating over a constant star-formation rate in a stellar population like in the Milky Way – about 10% of all massive stars are expected to be remnants of mergers from isolated binary evolution (Podsiadlowski et al. 1992; de Mink et al. 2014). Indeed, the binary evolution computations of Henneco et al. (2024) show that at least 16% of massive, mass-transferring binaries will merge, and this fraction goes up to if fully conservative mass transfer is assumed (Henneco 2025).
@@
Ever since the discovery of surface magnetic fields in the A-type star 78 Vir by Babcock (1947), the origin of strong surface magnetism in massive OBA stars is debated. It has long been speculated that strong and large-scale magnetic fields are produced in stellar mergers and that they can persist over the long evolutionary timescales of stars (see e.g. Ferrario et al. 2009; Wickramasinghe et al. 2014). Today, direct simulations of essentialy every flavor of stellar mergers and other dynamic stellar phases show the efficient amplification of magnetic fields. This is the case in mergers of binary white dwarfs (e.g. Zhu et al. 2015; Pakmor et al. 2024), white dwarfs and neutron stars (e.g. Morán-Fraile et al. 2024), binary neutron stars (e.g. Kiuchi et al. 2024), low- and high-mass main-sequence (MS) stars (e.g. Schneider et al. 2019, hereafter Paper I; Ryu et al. 2025; Vynatheya et al. 2025) and also common-envelope events (e.g. Ohlmann et al. 2016b; Ondratschek et al. 2022; Vetter et al. 2024; Gagnier & Pejcha 2024).
@@ 2.2 Hydrodynamics simulations
For the merger simulations, we need to create initial conditions that consist of two stars that are about to merge. For these single stars, we generate stable models by following the method outlined in Ohlmann et al. (2017). We compute the stellar structure with Mesa and map it to a grid with cell centers on spherical shells distributed according to a HEALPix tesselation. Then, we run a relaxation because slight mismatches between pressure gradients and gravitational forces can induce spurious velocities that need to be damped away. We carry out the relaxation over ten dynamical timescales as given by the sound crossing time to ensure that the models remain stable. We give some parameters of the two initial stellar models in Table 1.
@@
where is a timescale on which the components merge, and is the index of the cell. For this step, we take s and run the simulation for about 1.5 orbits with the loss term and continue without it from there on. For comparison, we start one model without the loss term after only half this time and it shows a very similar evolution. The simulations use 3.4 to 4 million cells in the high resolution runs and about 0.43 to 0.45 million cells in the low resolution runs. We set up the initial magnetic field as a dipole field, similar to that in Ohlmann et al. (2016b), with a surface field of 1 G. We run one additional simulation without magnetic fields for comparison. We provide an overview of the models in Table 2.
@@ 3.2 Evolution of the merger
The morphology of the flow during the merger is depicted in Fig. 2, showing slices of the density through the orbital plane and perpendicular to it, and in Fig. 3, showing slices of the passive scalar tracing material of the primary star through the orbital plane. Initially, mass transfer from the primary star to the secondary star sets in via the inner Lagrange point L1. This stream hits the secondary and leads to shear flows near the surface of the secondary. In addition, mass and also angular momentum are lost through the outer Lagrange points L2 and L3 (see, e.g., Fig. 2 at ), thereby shrinking the orbit. The orbital shrinkage accelerates until the stars merge: the initially more massive primary star is disrupted and distributed around the secondary star (see Figs. 2 and 3 at ). Directly after the merger, the cores of the stars are still separate (see central part of Fig. 3 at ). Shear in this central part causes hydrodynamical instabilities, mixing the cores and resulting in a symmetric merger product (Figs. 2 and 3 at ). Due to the large angular momentum of the binary, a disk-like structure forms around the rotating core with a mass of about (see Fig. 2 at in the – plane). The central core mainly originates from the less massive secondary, while the fraction of material from the primary increases outwards (Fig. 3). The post-merger structure is described in more detail in Sect. 3.4.
The simulations of the merger start after the orbit of the binary system was shrunk to (Model 1–3) or (Model 4), corresponding to orbital periods of about d and d, respectively (cf. Sect. 2.2). Strong mass transfer () develops from the more massive primary star onto the secondary. The orbital separation decreases until the stars merge after roughly 23 (Model 1) to 35 (Model 4) orbits111For larger initial separations, the system becomes even more stable and takes more orbits to merge, rendering the simulations computationally unfeasible. (Fig. 4). As expected, Model 4 with a larger initial separation takes longer to merge than Models 1–3 with a smaller initial separation. Here, we define as the point in time where the orbital shrinkage becomes large, i.e., where the corresponding time scale becomes small and fulfills (where is the semi-major axis). Fig. 4 also shows that the distance varies with the orbital motions, i.e., the orbit is slightly eccentric. This is probably caused by the system not being perfectly Keplerian after the first step of artificially shrinking the binary orbit. The eccentricity is smaller for the high-resolution Model 1.
@@ 3.3 Magnetic field amplification
The amplification continues in the mass stream from the primary to the secondary as it creates shear flows outside the Roche lobes of both stars (Fig. 6 at d). At this stage, we find that the flow is unstable to the magneto-rotational instability (MRI, Balbus & Hawley 1998). We estimate the length and time scale of the fastest growing MRI channel in our simulations (cf. Rembiasz et al. 2016): this yields a timescale of about 0.5 d and a wavelength of about which is similar to the scales in our simulation (time scale: 0.2–1 d, length scale:). Moreover, the strong amplification of the magnetic field starts as soon as the wavelength of the fastest growing MRI channel is resolved on the grid.
@@
The amplification is slowing down and reaches saturation at later times, when the ratio of magnetic to turbulent kinetic energy approaches the percent level and comes closer to equipartition (Fig. 5, middle panel). As a proxy for the turbulent kinetic energy, we use the kinetic energy in the radial and direction, , which is not affected by the large-scale angular motion (see the comparison in Fig. 5, bottom panel). We even find super equipartition, because some of the differential rotational energy from the large-scale ordered motion at is converted into magnetic energy.
The length scales can also be identified in power spectra of the kinetic and magnetic energy (Fig. 7). Most of the kinetic energy resides at large scales and low wave numbers for all times, as expected for the rotation of the stellar cores and of the post-merger product (Fig. 7, upper panel). The power spectrum follows a power law to smaller scales (exponent between and ), probably due to turbulence seeded by flow instabilities. The exponents are smaller in magnitude than the Kolmogorov exponent of that would be expected for a homogenous, isotropic turbulent cascade. Plots of power spectra of the magnetic energy (Fig. 7, lower panel), however, show that magnetic energy is injected at small scales (about –) and propagates to larger scales during the amplification (see d to d). After the cores have merged, redistribution of the magnetic energy leads to a similar spectrum at 5 d as for the kinetic energy, with most of the energy at large scales (about –). The exponents of the power law have been obtained by a fit to for between 0.1 and and are 1.38, , and for d, d, and d, respectively. This behavior is similar to the MRI simulations by Bhat et al. (2016), in which large-scale modes only grow late in the evolution when the amplification nears saturation. As mentioned there, large-scale magnetic fields seem to appear when the amplification process saturates (see also Ebrahimi et al. 2009). We suggest that this energy conversion from small to large scales is partially due to the large-scale ordered fluid motions winding up magnetic fields into a dominant toroidal component and an effective, large-scale dynamo that also explains the magnetic amplification beyond equipartition with the turbulent energy, akin to the case of WD mergers studied in Pakmor et al. (2024).
@@ 3.4 Structure of the merger product
The large angular momentum in the binary system leads to the formation of a disk around a rotating core in the merger product (compare Fig. 2 at d). The density profile shows strong asymmetries between polar and equatorial regions, indicating that the boundary between core and disk is located roughly at (Fig. 8). Moreover, the mass profile increases only in the equatorial region for radii above (Fig. 8, lower panel). Thus, almost all mass for radii above is contained in the equatorial region, i.e., in a disk-like structure. At d, the core contains a mass of about , and the disk a mass of about . Moreover, the disk comprises 58% of the total angular momentum.
@@
The velocity at the core–disk boundary is about 700 km s-1 at d, i.e. 76% of the critical Keplerian rotational velocity. As can be seen in Fig. 9 (second panel) the rotation curve in the disk always stays sub-Keplerian. After six days, this disk has completed about 170 rotations at the base and about 0.3 rotations at the edge. The ratio of kinetic energy to gravitational energy decreases from 0.07 at d to 0.05 at d and stays well below the limit of 0.14 for non-axisymmetric instabilities in rotating fluids (Ostriker & Tassoul 1969; Chandrasekhar 1970; Friedman & Schutz 1975; Lai & Shapiro 1995). Thus, no secular or dynamical, non-axisymmetric instabilities of the configuration are expected.
The magnetic field strength (Fig. 9) still increases throughout the merger product at the end of the simulation, which means that the amplification is not yet completely saturated. Thus, there still seems to be an active slow dynamo, probably connected to the large-scale rotation of the merger product. Ultimately, the simulation would need to be continued to follow the amplification until it is completely saturated.
@@
As already seen during the evolution of the merger (cf. Fig. 3), the very central part of the merger product originates from the secondary star, whereas the primary star is disrupted around the core. Only the inner of the product are dominated by material from the secondary; the rest of the core stems to similar parts from both stars. The disk, however, is predominantly formed from material of the primary star. This distribution is expected from the entropy and thus buoyancy profiles of the progenitor stars: due to the similar mass and hence also similar evolutionary stage, the entropy is similar and mixing is expected over large parts of the profile. Because the secondary has a lower mass, its entropy is slightly smaller and thus its material dominates the central part. This behaviour is similar to the case of direct collisions as explored by Glebbeek & Pols (2008) and Glebbeek et al. (2013), especially their case S.
@@
The magnetic field structure of the final merger product is visualized in Fig. 10. The field structure is mostly toroidal; the energy of the toroidal magnetic field component is about 80–85% of the total magnetic energy. In this range, magnetic field configurations are expected to be stable in stars (Braithwaite & Spruit 2004; Braithwaite & Nordlund 2006a). Moreover, the configuration remains stable after the final merger, and this time already corresponds to a few Alfvén timescales. However, in the regime of fast rotation (), the evolutionary timescale of the magnetic field is , where is the Alfvén timescale (Braithwaite & Cantiello 2013). Hence, significant changes in the magnetic field are not expected to be observed over the timescales simulated here after the field has reached an equilibrium configuration.
Due to Ohmic resistivity, the magnetic field is expected to dissipate and the corresponding time scale depends on the square of the coherence scale of the magnetic field. Using Spitzer resistivity, we can estimate the dissipation time scale as
where is the temperature of the plasma. In Fig. 11, we plot the azimuthal component of the magnetic field which shows that the magnetic field is coherent on length scales of roughly . With temperatures between and , we get timescales between and , larger than the lifetime of the star, which should rather be about . Hence, the magnetic field of the star should remain present at least until it reaches the main sequence. Modeling the subsequent stellar evolution in more detail shows that the merger product could evolve into an MS star with strong magnetic fields, similar to Sco (Paper I; Schneider et al. 2020).